If we put a 12 kg piece of Gold with a temperature of 125 degrees Celsius into a bucket of water with a mass of 8 kg and a temperature of 2 degrees Celsius. What would the final temperature be? *



1
Expert's answer
2021-04-20T16:48:58-0400

Solution.

"m_g=12kg;"

"T_g=125^oC;"

"c_g=125J\/(kgK);"

"m_w=8kg;"

"t_w=2^oC;"

"c_w=4200J\/(kgK);"

"T_f-?;"

"c_gm_g(T_g-T_f)=c_wm_w(T_f-T_w);"

"T_f=\dfrac{c_gm_gT_g+c_wm_wT_w}{c_gm_g+c_wm_w};"

"T_f=\dfrac{125J/(kgK)\sdot12kg\sdot125^oC+4200J/(kgK)\sdot8kg\sdot2^oC}{125J/(kgK)\sdot12kg+4200J/(kgK)\sdot8kg}=7^oC;"

Answer: "T_f=7^oC."